# Confluent functions, Laguerre polynomials and their (generalized) bilinear integrals

### v1

@@ Abstract
1 Introduction
@@
${\cal F}_{\alpha}$ $\coloneqq-z\partial_{z}^{2}-(1+\alpha-z)\partial_{z}.$ (1.1)
Kummer’s confluent function $F_{\theta,\alpha}(z)\coloneqq{}_{1}F_{1}(a;c;z)$ is an eigenfunction of ${\cal F}_{\alpha}$ with eigenvalue $\frac{1+\theta+\alpha}{2}$, where following [4, 5] instead of $a,c$ we prefer to use the parameters
$\alpha\coloneqq c-1,\qquad\theta:=-c+2a.$ (1.2)
Besides Kummer’s function, (1.1) possesses a second type of distinguished eigenfunction with eigenvalues $\frac{1+\theta+\alpha}{2}$, called Tricomi’s confluent functions, which we denote $U(a;c;z)$ or, in the parameters (1.2), $U_{\theta,\alpha}(z)$. Tricomi’s functions usually have a better behavior near infinity than Kummer’s functions.
@@
$\quad\int_{0}^{\infty}U_{\theta_{1},\alpha}(z)U_{\theta_{2},\alpha}(z){\rm e}^{-z}z^{\alpha}{\rm d}z.$ (1.3)
@@
It is known that for $|\mathrm{Re}\alpha|<1$ (1.3) can be computed explicitly, see e.g. [6]. For $|\mathrm{Re}\alpha|\geq 1$, (1.3) is divergent — except for very specific combinations of the parameters — because of the behavior of the integrand near 00. However, when one applies the so-called generalized integral, one can compute (1.3) for all values of $\alpha$. For $\alpha\in{\mathbb{Z}}$, $\alpha\neq 0$ one needs to apply the anomalous generalized integral, and one obtains a much more complicated expression.
@@
For special values of the parameters, Tricomi’s functions coincide up to a coefficient with the well-known Laguerre polynomials. Their classical definition uses the so-called Rodriguez formula
$L_{n}^{\alpha}(z)\coloneqq\frac{1}{n!}{\rm e}^{z}z^{-\alpha}\partial_{z}^{n}{\rm e}^{-z}z^{n+\alpha},\quad n\in{\mathbb{N}}_{0},$ (1.4)
@@
and applies to all $\alpha\in{\mathbb{C}}$. They are, however, orthogonal with respect to the positive definite measure ${\mathrm{d}}\mu(z)=z^{\alpha}\mathrm{e}^{-z}{\mathrm{d}}z$ on $L^{2}[0,\infty[$ only if $\alpha>-1$.
The orthogonality relations for Laguerre polynomials, valid for $\mathrm{Re}\alpha>-1$,
$\int_{0}^{\infty}L_{m}^{\alpha}(z)L_{n}^{\alpha}(z)z^{\alpha}{\rm e}^{-z}{\mathrm{d}}z=\frac{\Gamma(1+n+\alpha)}{n!}\delta_{m,n},$ (1.5)
@@
belong to the standard knowledge. Again, using the generalized integral, we extend the relations (1.5) to all $\alpha\in{\mathbb{C}}$. For $-\alpha\not\in{\mathbb{N}}$ we still have the same expression as in (1.5). In particular, $L_{n}^{\alpha}$ are (pseudo-)orthogonal to one another for $m\neq n$. The situation is more complicated for $-\alpha\in{\mathbb{N}}$, where the generalized integral becomes anomalous.
@@
One can derive the bilinear integrals of Laguerre polynomials from those of Tricomi’s functions. However, we prefer a direct method, which mimics the usual proof of the orthogonality relations based on the Rodriguez formula and repeated integration by parts, known from standard textbooks. Interestingly, integration by parts works quite well also for generalized integrals. This is not obvious, so we dedicate a subsection in Appendix A to the description of integration by parts in the context of generalized integrals.
@@
For $\alpha\not\in[-1,\infty[$ (typically real, preferably integer), Laguerre polynomials are of interest in several fields of mathematics, see for example [24, 25, 26, 14]. Apparently, the value $\alpha=-1$ is especially interesting in applications [2].
@@
In a recent work [7], some of us (with collaborators) used the generalized integral, a concept that goes back to independent considerations of Riesz [23] and Hadamard [9, 10], to give a meaning to integrals over classically non-integrable functions. The generalized integral is a linear continuation of the standard integral. It is closely related to the extension of homogeneous distributions [11, 22] and the so-called barred integral [18]. We find it interesting that replacing the integral by the generalized integral, we are able to compute bilinear integrals of the form (1.3) and (1.5) for arbitrary complex $\alpha$.
@@
Our paper is not the first to address generalized orthogonality relations of the classical orthogonal polynomials. There is some literature on the subject [12, 15, 16, 17, 19, 13], where generalized orthogonality relations for some $\alpha\in{\mathbb{C}}$ with $\mathrm{Re}(\alpha)<-1$ are often described by changing the integral measure, typically allowing for distributional measures. The generalized integral is a natural tool that allows us to treat all $\alpha$ simultaneously. We have not seen such a comprehensive description in the literature.
@@
Here is the plan of our paper. Section 2 is devoted to the confluent equation for general parameters. First we recall its basic theory, following mostly [4, 5]. In particular, we review the definitions of four standard solutions of the confluent equation. Three of them typically blow up at infinity, and only one of them, the Tricomi function, can be used in bilinear (generalized) integrals of the form (1.3). The main new result of this section are the expressions for these integrals for all parameters.
@@
Section 3 is devoted to Laguerre polynomials. Again, we start with a concise exposition of their theory. Then we compute their bilinear integrals. We use two methods. Our first method is based on integration by parts, which interestingly works well also in the case of generalized integrals. Then we give an alternative derivation: we show how bilinear integrals of Laguerre polynomials can be obtained as special cases of bilinear integrals of Tricomi functions obtained in Section 2.
@@
2 Confluent equation
2.1 Hypergeometric ${}_{1}F_{1}$ function
@@
${}_{1}F_{1}(a;c;z)$ $\coloneqq\sum_{n=0}^{\infty}\frac{(a)_{n}}{(c)_{n}}\frac{z^{n}}{n!}.$ (2.1)
@@
${}_{1}{\bf F}_{1}(a;c;z)$ $\coloneqq\frac{{}_{1}F_{1}(a;c;z)}{\Gamma(c)}=\sum_{n=0}^{\infty}\frac{(a)_{n}}{\Gamma(c+n)}\frac{z^{n}}{n!}.$ (2.5)
@@
$\frac{1}{2\pi{\rm i}}\int\limits_{]-\infty,(0,z)^{+},-\infty[}t^{a-c}{\rm e}^{t}(t-z)^{-a}{\rm d}t$ $=$ ${}_{1}{\bf F}_{1}(a;c;z),$ (2.6)
for $\mathrm{Re}a>0,\ \mathrm{Re}(c-a)>0$,
$\frac{1}{\Gamma(a)\Gamma(c-a)}\int\limits_{[1,+\infty[}{\rm e}^{\frac{z}{t}}t^{-c}(t-1)^{c-a-1}{\rm d}t$ $=$ ${}_{1}{\bf F}_{1}(a;c;z),$ (2.7)
and for $\mathrm{Re}(c-a)>0$ and $a\notin{\mathbb{N}}$,
$\frac{\Gamma(1-a)}{2\pi{\rm i}\Gamma(c-a)}\int\limits_{[1,0^{+},1]}{\rm e}^{\frac{z}{t}}(-t)^{-c}(-t+1)^{c-a-1}{\rm d}t$ $=$ ${}_{1}{\bf F}_{1}(a;c;z).$ (2.8)
The meaning of the contours in (2.6) and (2.8) is the following. In (2.6), we start at $\infty$- ∞, go around both 00 and $z$ in the positive sense and go back to $\infty$- ∞. Similarly, in (2.8), we start from $1$, go around 00 in the positive sense and go back to $1$.
@@
${}_{1}F_{1}(c+m;c;z)$ $=\frac{z^{1-c}}{(c)_{m}}\partial_{z}^{m}z^{c-1+m}{\rm e}^{z},$ (2.9)
${}_{1}{\bf F}_{1}(c+m;c;z)$ $=\frac{z^{1-c}}{\Gamma(c+m)}\partial_{z}^{m}z^{c-1+m}{\rm e}^{z}.$ (2.10)
@@
$\partial_{z}^{-1}f(z)\coloneqq\mathrm{gen}\int_{0}^{z}f(y){\rm d}y,$ (2.11)
provided that (2.11) is non-anomalous and that the function is integrable in the generalized sense often enough.
2.2 Hypergeometric ${}_{2}F_{0}$ function
@@
where $|\arg c-\pi|<\pi-\epsilon$, $\epsilon>0$. It extends to an analytic function on the universal cover of \$\0{\mathbb{C}}\backslash\{0\}$ with a branch point of an infinite order at 0. It has the following asymptotic expansion:
@@
We have an integral representation for $\mathrm{Re}a>0$,
@@
2.3 Equivalence between the ${}_{1}F_{1}$ and ${}_{2}F_{0}$ equations
@@
2.4 Confluent equation in Lie-algebraic parameters
@@
$\begin{array}[]{lll}\alpha\coloneqq c-1=a-b,&&\theta:=-c+2a=-1+a+b;\\[8.61108pt] a=\frac{1+\alpha+\theta}{2},&b=\frac{1-\alpha+\theta}{2},&c=1+\alpha.\end{array}$
@@
$\Big{(}z\partial_{z}^{2}+(1+\alpha-z)\partial_{z}-\frac{1}{2}(1+\theta+\alpha)\Big{)}f(z)$ $=0,$ (2.13)
$\Big{(}w^{2}\partial_{w}^{2}+\big{(}-1+(2+\theta)w\big{)}\partial_{w}+\big{(}\tfrac{1+\theta}{2}\big{)}^{2}-\big{(}\tfrac{\alpha}{2}\big{)}^{2}\Big{)}f(w)$ $=0.$ (2.14)
@@
${\cal F}_{\alpha}(z,\partial_{z})$ $\coloneqq-z\partial_{z}^{2}-(1+\alpha-z)\partial_{z}$ (2.15)
$=-z^{-\alpha}{\rm e}^{z}\partial_{z}z^{1+\alpha}{\rm e}^{-z}\partial_{z},$ (2.16)
~$\tilde{\cal F}_{\theta}(z,\partial_{w})$ $\coloneqq-w^{2}\partial_{w}^{2}-\big{(}-1+(2+\theta)w\big{)}\partial_{w}$ (2.17)
$=-w^{-\theta}{\rm e}^{-w^{-1}}\partial_{w}w^{2+\theta}{\rm e}^{w^{-1}}\partial_{w}.$ (2.18)
@@
$\Big{(}-{\cal F}_{\alpha}(z,\partial_{z})-\frac{1}{2}(1+\theta+\alpha)\Big{)}f(z)$ $=0,$ (2.19)
(~$\Big{(}-\tilde{\cal F}_{\theta}(w,\partial_{w})+\big{(}\tfrac{1+\theta}{2}\big{)}^{2}-\big{(}\tfrac{\alpha}{2}\big{)}^{2}\Big{)}f(w)$ $=0.$ (2.20)
@@
$F_{\theta,\alpha}(z)$ $\coloneqq{}_{1}{F}_{1}\Bigl{(}\frac{1+\alpha+\theta}{2};1+\alpha;z\Bigr{)},$ (2.21)
${\bf F}_{\theta,\alpha}(z)$ $\coloneqq{}_{1}{\bf F}_{1}\Bigl{(}\frac{1+\alpha+\theta}{2};1+\alpha;z\Bigr{)}=\frac{1}{\Gamma(\alpha+1)}F_{\theta,\alpha}(z),$ (2.22)
$U_{\theta,\alpha}(z)$ $\coloneqq U\bigl{(}\tfrac{1+\alpha+\theta}{2};1+\alpha;z\bigr{)}=z^{\frac{-1-\alpha-\theta}{2}}{{}_{2}F_{0}}\bigl{(}\tfrac{1+\alpha+\theta}{2},\tfrac{1-\alpha+\theta}{2};-;-z^{-1}\bigr{)}.$ (2.23)
@@
$F_{\theta,\alpha}(z)$ $={\rm e}^{z}F_{-\theta,\alpha}(-z),$ $\sim 1\text{ at }0;$ (2.24)
$z^{-\alpha}F_{\theta,-\alpha}(z)$ $=z^{-\alpha}{\rm e}^{z}F_{-\theta,-\alpha}(-z),$ $\sim z^{-\alpha}\text{ at }0;$ (2.25)
$U_{\theta,\alpha}(z)$ $=z^{-\alpha}U_{\theta,-\alpha}(z),$ $\sim z^{-a}\text{ at }+\infty;$ (2.26)
${\rm e}^{z}U_{-\theta,\alpha}(-z)$ $=(-z)^{-\alpha}{\rm e}^{z}U_{-\theta,-\alpha}(-z),$ $\sim(-z)^{b-1}{\rm e}^{z}\text{ at }-\infty.$ (2.27)
2.5 Connection formulas
The space of solutions of the confluent equation is 2-dimensional. Therefore, the Tricomi function for $\alpha\not\in{\mathbb{Z}}$ can be expressed in terms of Kummer’s functions:
$U_{\theta,\alpha}(z)$ $=\frac{\pi}{\sin\pi\alpha}\Bigg{(}-\frac{{\bf F}_{\theta,\alpha}(z)}{\Gamma\left(\frac{1+\theta-\alpha}{2}\right)}+\frac{z^{-\alpha}{\bf F}_{\theta,-\alpha}(z)}{\Gamma\left(\frac{1+\theta+\alpha}{2}\right)}\Bigg{)},$ (2.28)
${\rm e}^{z}U_{-\theta,\alpha}(-z)$ $=\frac{\pi}{\sin\pi\alpha}\Bigg{(}-\frac{{\bf F}_{\theta,\alpha}(z)}{\Gamma\left(\frac{1-\theta-\alpha}{2}\right)}+\frac{(-z)^{-\alpha}{\bf F}_{\theta,-\alpha}(z)}{\Gamma\left(\frac{1-\theta+\alpha}{2}\right)}\Bigg{)}.$ (2.29)
2.6 Degenerate case

### v2

@@ Abstract
Keywords: special functions, confluent equation, orthogonal polynomials, regularization.
1. Introduction
@@
${\cal F}_{\alpha}$ $\coloneqq-z\partial_{z}^{2}-(1+\alpha-z)\partial_{z}.$ (1)
Kummer’s confluent function $F_{\theta,\alpha}(z)\coloneqq{}_{1}F_{1}(a;c;z)$ is an eigenfunction of ${\cal F}_{\alpha}$ with eigenvalue $\frac{1+\theta+\alpha}{2}$, where following [1, 2, 3] instead of $a,c$ we prefer to use the parameters
$\alpha\coloneqq c-1,\qquad\theta:=-c+2a.$ (2)
Besides Kummer’s function, (1) possesses a second type of distinguished eigenfunction with eigenvalues $\frac{1+\theta+\alpha}{2}$, called Tricomi’s confluent functions, which we denote $U(a;c;z)$ or, in the parameters (2), $U_{\theta,\alpha}(z)$. Tricomi’s functions usually have a better behavior near infinity than Kummer’s functions.
@@
$\quad\int_{0}^{\infty}U_{\theta_{1},\alpha}(z)U_{\theta_{2},\alpha}(z){\rm e}^{-z}z^{\alpha}{\rm d}z.$ (3)
@@
It is known that for $|\Re(\alpha)|<1$ (3) can be computed explicitly, see e.g. [4]. For $|\Re(\alpha)|\geq 1$, (3) is divergent — except for very specific combinations of the parameters — because of the behavior of the integrand near 0. However, when one applies the so-called generalized integral, one can compute (3) for all values of $\alpha$. For $\alpha\in{\mathbb{Z}}$, $\alpha\neq 0$, one needs to apply the anomalous generalized integral, and one obtains a much more complicated expression.
@@
For special values of the parameters, Tricomi’s functions coincide up to a coefficient with the well-known Laguerre polynomials. Their classical definition uses the so-called Rodrigues formula
$L_{n}^{\alpha}(z)\coloneqq\frac{1}{n!}{\rm e}^{z}z^{-\alpha}\partial_{z}^{n}{\rm e}^{-z}z^{n+\alpha},\quad n\in{\mathbb{N}}_{0},$ (4)
@@
and applies to all $\alpha\in{\mathbb{C}}$. They are, however, orthogonal with respect to the positive definite measure ${\mathrm{d}}\mu(z)=z^{\alpha}{\rm e}^{-z}{\mathrm{d}}z$ on $L^{2}[0,\infty[$ only if $\alpha>-1$.
The orthogonality relations for Laguerre polynomials, valid for $\Re(\alpha)>-1$,
$\int_{0}^{\infty}L_{m}^{\alpha}(z)L_{n}^{\alpha}(z)z^{\alpha}{\rm e}^{-z}{\mathrm{d}}z=\frac{\Gamma(1+n+\alpha)}{n!}\delta_{m,n},$ (5)
@@
belong to the standard knowledge. Again, using the generalized integral, we extend the relations (5) to all $\alpha\in{\mathbb{C}}$. For $\alpha\not\in{\mathbb{N}}$ we still have the same expression as in (5). In particular, $L_{n}^{\alpha}$ are (pseudo-)orthogonal to one another for $m\neq n$. The situation is more complicated for $\alpha\in{\mathbb{N}}$, where the generalized integral becomes anomalous.
@@
One can derive the bilinear integrals of Laguerre polynomials from those of Tricomi’s functions. However, we prefer a direct method, which mimics the usual proof of the orthogonality relations based on the Rodrigues formula and repeated integration by parts, known from standard textbooks. Interestingly, integration by parts works quite well also for generalized integrals. This is not obvious, so we dedicate a subsection in Appendix A to the description of integration by parts in the context of generalized integrals.
@@
For $\alpha\not\in[-1,\infty[$ (typically real, preferably integer), Laguerre (or other orthogonal) polynomials are of interest in several fields of mathematics, see for example [5, 6, 7, 8, 9]. Apparently, the value $\alpha=-1$ is especially interesting in applications [10].
@@
In a recent work [11], some of us (with collaborators) used the generalized integral, a concept that goes back to independent considerations of Riesz [12] and Hadamard [13, 14], to give a meaning to integrals over classically non-integrable functions. The generalized integrals of these functions appear as higher-dimensional generalizations of Green’s functions for Laplacians perturbed by point interactions [15], forming toy models for various facets of renormalization in quantum field theory. The generalized integral is a linear continuation of the standard integral. It is closely related to the extension of homogeneous distributions [16, 17] and the so-called barred integral [18]. We find it interesting that replacing the integral by the generalized integral, we are able to compute bilinear integrals of the form (3) and (5) for arbitrary complex $\alpha$.
@@
Our paper is not the first to address generalized orthogonality relations of the classical orthogonal polynomials. There is some literature on the subject [19, 20, 21, 22, 23, 24, 25], where generalized orthogonality relations for some $\alpha\in{\mathbb{C}}$ with $\Re(\alpha)<-1$ are often described by changing the integral measure, typically allowing for distributional measures. The generalized integral is a natural tool that allows us to treat all $\alpha$ simultaneously. We have not seen such a comprehensive description in the literature.
Integrals of expressions containing confluent functions can be found in many sources. However, most of them involve a single confluent function, such as those in [26], where they are derived with the so-called umbral technique. Integrals of products of two confluent functions are computed e.g. in [4]; they are less frequent in the literature. The identity (123), whose proof involves a telescoping series, belongs to a class of identities treated by umbral methods in [27].
@@
Here is the plan of our paper. Section 2. is devoted to the confluent equation for general parameters. First we recall its basic theory, following mostly [1, 2]. In particular, we review the definitions of four standard solutions of the confluent equation. Three of them typically blow up at infinity, and only one of them, the Tricomi function, can be used in bilinear (generalized) integrals of the form (3). The main new result of this section are the expressions for these integrals for all parameters.
@@
Section 3. is devoted to Laguerre polynomials. Again, we start with a concise exposition of their theory. Then we compute their bilinear integrals. We use two methods. Our first method is based on integration by parts, which interestingly works well also in the case of generalized integrals. Then we give an alternative derivation: we show how bilinear integrals of Laguerre polynomials can be obtained as special cases of bilinear integrals of Tricomi functions obtained in Section 2..
@@
2. Confluent equation
In this section we start with an overview of the confluent hypergeometric equation and its standard solutions. All the identities of this section until Subsection 2.7. can be found in one form or another in standard sources, such as [28]. Nevertheless, we believe that our presentation, influenced by [1, 2, 3], is quite different from the usual one. We stress that the confluent equation has two different but equivalent forms: it can be understood as an eigenvalue equation of the $_{1}F_{1}$ operator (27) and of the $_{2}F_{0}$ operator (28). These two forms are related to the so-called Lie-algebraic parameters (20), which are helpful in organizing the properties of the confluent equation.
For convenience, and as we employ them heavily, we recall the definition of the Pochhammer symbol and the harmonic numbers
$(a)_{n}=a(a+1)...(a+n-1),\quad H_{k}(a)=\sum_{j=0}^{k-1}\frac{1}{a+j},$ (6)
which admit a useful extension to complex parameters in terms of the Gamma and Digamma functions,
$(a)_{z}=\frac{\Gamma(a+z)}{\Gamma(a)},\quad H_{z}(a)=\psi(a+z)-\psi(a),\quad z\in{\mathbb{C}}\setminus(-a-{\mathbb{N}}_{0}).$ (7)
2.1. Hypergeometric $_{1}F_{1}$ function
@@
${}_{1}F_{1}(a;c;z)$ $\coloneqq\sum_{n=0}^{\infty}\frac{(a)_{n}}{(c)_{n}}\frac{z^{n}}{n!}.$ (8)
@@
${}_{1}{\bf F}_{1}(a;c;z)$ $\coloneqq\frac{{}_{1}F_{1}(a;c;z)}{\Gamma(c)}=\sum_{n=0}^{\infty}\frac{(a)_{n}}{\Gamma(c+n)}\frac{z^{n}}{n!}.$ (12)
@@
$\frac{1}{2\pi{\rm i}}\int\limits_{]-\infty,(0,z)^{+},-\infty[}t^{a-c}{\rm e}^{t}(t-z)^{-a}{\rm d}t$ $=$ ${}_{1}{\bf F}_{1}(a;c;z),$ (13)
for $\Re(a)>0,\ \Re(c-a)>0$,
$\frac{1}{\Gamma(a)\Gamma(c-a)}\int\limits_{[1,+\infty[}{\rm e}^{\frac{z}{t}}t^{-c}(t-1)^{c-a-1}{\rm d}t$ $=$ ${}_{1}{\bf F}_{1}(a;c;z),$ (14)
and for $\Re(c-a)>0$ and $a\notin{\mathbb{N}}$,
$\frac{\Gamma(1-a)}{2\pi{\rm i}\Gamma(c-a)}\int\limits_{[1,0^{+},1]}{\rm e}^{\frac{z}{t}}(-t)^{-c}(-t+1)^{c-a-1}{\rm d}t$ $=$ ${}_{1}{\bf F}_{1}(a;c;z).$ (15)
The meaning of the contours in (13) and (15) is the following. In (13), we start at -$\infty$, go around both 0 and $z$ in the positive sense and go back to -$\infty$. Similarly, in (15), we start from $1$, go around 0 in the positive sense and go back to $1$.
@@
${}_{1}F_{1}(c+m;c;z)$ $=\frac{z^{1-c}}{(c)_{m}}\partial_{z}^{m}z^{c-1+m}{\rm e}^{z},$ (16)
${}_{1}{\bf F}_{1}(c+m;c;z)$ $=\frac{z^{1-c}}{\Gamma(c+m)}\partial_{z}^{m}z^{c-1+m}{\rm e}^{z}.$ (17)
@@
$\partial_{z}^{-1}f(z)\coloneqq\mathrm{gen}\int_{0}^{z}f(y){\rm d}y,$ (18)
provided that (18) is non-anomalous and that the function is integrable in the generalized sense often enough.
2.2. Hypergeometric $_{2}F_{0}$ function
@@
where $|\arg c-\pi|<\pi-\epsilon$, $\epsilon>0$. It extends to an analytic function on the universal cover of $\{$${\mathbb{C}}\backslash\{0\}$ with a branch point of an infinite order at 0. It has the following asymptotic expansion:
@@
We have an integral representation for $\Re(a)>0$,
@@
2.3. Equivalence between the $_{1}F_{1}$ and $_{2}F_{0}$ equations
@@
2.4. Confluent equation in Lie-algebraic parameters
@@
$\begin{array}[]{lll}\alpha\coloneqq c-1=a-b,&&\theta:=-c+2a=-1+a+b;\\[8.61108pt] a=\frac{1+\alpha+\theta}{2},&b=\frac{1-\alpha+\theta}{2},&c=1+\alpha.\end{array}$ (20)
@@
$\Big(z\partial_{z}^{2}+(1+\alpha-z)\partial_{z}-\frac{1}{2}(1+\theta+\alpha)\Big)f(z)$ $=0,$ (21)
$\Big(w^{2}\partial_{w}^{2}+\big(-1+(2+\theta)w\big)\partial_{w}+\big(\tfrac{1+\theta}{2}\big)^{2}-\big(\tfrac{\alpha}{2}\big)^{2}\Big)f(w)$ $=0.$ (22)
@@
${\cal F}_{\alpha}(z,\partial_{z})$ $\coloneqq-z\partial_{z}^{2}-(1+\alpha-z)\partial_{z}$ (23)
$=-z^{-\alpha}{\rm e}^{z}\partial_{z}z^{1+\alpha}{\rm e}^{-z}\partial_{z},$ (24)
$\tilde{\cal F}_{\theta}(z,\partial_{w})$ $\coloneqq-w^{2}\partial_{w}^{2}-\big(-1+(2+\theta)w\big)\partial_{w}$ (25)
$=-w^{-\theta}{\rm e}^{-w^{-1}}\partial_{w}w^{2+\theta}{\rm e}^{w^{-1}}\partial_{w}.$ (26)
@@
$\Big(-{\cal F}_{\alpha}(z,\partial_{z})-\frac{1}{2}(1+\theta+\alpha)\Big)f(z)$ $=0,$ (27)
$\Big(-\tilde{\cal F}_{\theta}(w,\partial_{w})+\big(\tfrac{1+\theta}{2}\big)^{2}-\big(\tfrac{\alpha}{2}\big)^{2}\Big)f(w)$ $=0.$ (28)
@@
$F_{\theta,\alpha}(z)$ $\coloneqq{}_{1}{F}_{1}\Bigl(\frac{1+\alpha+\theta}{2};1+\alpha;z\Bigr),$ (29)
${\bf F}_{\theta,\alpha}(z)$ $\coloneqq{}_{1}{\bf F}_{1}\Bigl(\frac{1+\alpha+\theta}{2};1+\alpha;z\Bigr)=\frac{1}{\Gamma(\alpha+1)}F_{\theta,\alpha}(z),$ (30)
$U_{\theta,\alpha}(z)$ $\coloneqq U\bigl(\tfrac{1+\alpha+\theta}{2};1+\alpha;z\bigr)=z^{\frac{-1-\alpha-\theta}{2}}{{}_{2}F_{0}}\bigl(\tfrac{1+\alpha+\theta}{2},\tfrac{1-\alpha+\theta}{2};-;-z^{-1}\bigr).$ (31)
@@
$F_{\theta,\alpha}(z)$ $={\rm e}^{z}F_{-\theta,\alpha}(-z),$ $\sim 1\text{ at }0;$ (32)
$z^{-\alpha}F_{\theta,-\alpha}(z)$ $=z^{-\alpha}{\rm e}^{z}F_{-\theta,-\alpha}(-z),$ $\sim z^{-\alpha}\text{ at }0;$ (33)
[…]
2.5. Connection formulas
The space of solutions of the confluent equation is 2-dimensional. Therefore, the Tricomi function for $\alpha\not\in{\mathbb{Z}}$ can be expressed in terms of Kummer’s functions via the following connection formulas, whose derivation can be found in [28] and [2]:
$U_{\theta,\alpha}(z)$ $=\frac{\pi}{\sin\pi\alpha}\Bigg(-\frac{{\bf F}_{\theta,\alpha}(z)}{\Gamma\left(\frac{1+\theta-\alpha}{2}\right)}+\frac{z^{-\alpha}{\bf F}_{\theta,-\alpha}(z)}{\Gamma\left(\frac{1+\theta+\alpha}{2}\right)}\Bigg),$ (36)
${\rm e}^{z}U_{-\theta,\alpha}(-z)$ $=\frac{\pi}{\sin\pi\alpha}\Bigg(-\frac{{\bf F}_{\theta,\alpha}(z)}{\Gamma\left(\frac{1-\theta-\alpha}{2}\right)}+\frac{(-z)^{-\alpha}{\bf F}_{\theta,-\alpha}(z)}{\Gamma\left(\frac{1-\theta+\alpha}{2}\right)}\Bigg).$ (37)
2.6. Degenerate case